OpenStudy (anonymous):

A company has started selling a new type of smartphone at the price of $110-0.1x where x is the number of smartphones manufactured per day. the parts for each phone cost $50 and labor and overhead for running the plant cost $4000 per day. How many smartphones should the company manufacture and sell per day to maximize profit?

OpenStudy (anonymous):

\[ x=\text{number of phones per day}\\ \text{selling price per day:}\quad p(x)=110-0.1x\\ \text{cost price per day:}\quad c(x) = 50x+4000\\ \;\\ {\rm profit:}\quad P(x)=p(x)-c(x) \]

OpenStudy (anonymous):

so, what will be the function for profit?

OpenStudy (anonymous):

(110-.1x) -(50x+4000) but i have to take the derivative of it b/c i need to show a sine chart

OpenStudy (anonymous):

you have to take the derivative: correct why: because, the question is asking to find the "x" for which P(x) is maximum

OpenStudy (anonymous):

but there is a small mistake I made here. reading the question again, 110-0.1x id the cost of one phone

OpenStudy (anonymous):

so multiply the whole cost by x?

OpenStudy (anonymous):

so, the correct selling price: \[p(x)=x(110-0.1x)=110x-0.1x^2\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oh... okay thank you so much :)

OpenStudy (anonymous):

and you proceed with the first derivative

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what did you get for the critical point?

OpenStudy (anonymous):

o'course, in our analysis, we are assuming that all the phones that are manufactured per day will be sold in one day

OpenStudy (anonymous):

yes and i got 300 as the critical number

OpenStudy (anonymous):

good. you can now check that P(x) is indeed maximum at x=300 using the second derivative test.

OpenStudy (anonymous):

what's the second derivative test?

OpenStudy (anonymous):

can't you also check it using a sine chart

OpenStudy (anonymous):

you find the second derivative of P(x) and for each critical point, you check the sign of the second derivative. what is a sine chart?

OpenStudy (anonymous):

its a number line and you plug in the critical number(s) into the function. Then you plug in a number less than your critical number and one greater than the critical number into the function. determine whether the F(x) is greater than or less than that of your F(critical#)

OpenStudy (anonymous):

yes. that is the manual version of second derivative. second derivative tells you the same exact thing without the need to picturize it and in numbers.

OpenStudy (anonymous):

oh i like the visual better but that's just my preference

OpenStudy (anonymous):

yep. so there you go then.

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

you are welcome